# Distinguishing Predictive and Generative AI in Regulation

### v2

@@ Abstract.
Over the past decade, policymakers have developed a set of regulatory tools to ensure AI development aligns with key societal goals. Many of these tools were initially developed in response to concerns with predictive AI and therefore encode certain assumptions about the nature of AI systems and the utility of certain regulatory approaches. With the advent of generative AI, however, some of these assumptions no longer hold, even as policymakers attempt to maintain a single regulatory target that covers both types of AI.
In this paper, we identify four distinct aspects of generative AI that call for meaningfully different policy responses. These are the generality and adaptability of generative AI that make it a poor regulatory target, the difficulty of designing effective evaluations, new legal concerns that change the ecosystem of stakeholders and sources of expertise, and the distributed structure of the generative AI value chain.
In light of these distinctions, policymakers will need to evaluate where the past decade of policy work remains relevant and where new policies, designed to address the unique risks posed by generative AI, are necessary. We outline three recommendations for policymakers to more effectively identify regulatory targets and leverage constraints across the broader ecosystem to govern generative AI.
isbn:
@@ 1. Introduction
Recent advances in AI have drawn intense interest from policymakers. In the past year alone, policymakers have proposed and enacted a litany of governance measures, from the European Union AI Act to a growing number of state and local initiatives in the United States (European Parliament, 2024; Curry, 2024). These efforts reflect divergent priorities in balancing innovation with oversight.
Some treat the recent advent of generative AI as a fundamentally new class of technology whose scale and autonomy distinguish it from others (Bostrom, 2014; Altman et al., 2023). In this view, generative AI introduces distinct risks and capabilities that existing regulatory regimes are ill-equipped to handle (Anderljung et al., 2023). As a result, proponents argue that addressing these challenges requires entirely new laws and institutional structures (Nichols, 2023; Bureau of Industry and Security, U.S. Department of Commerce, 2025). Some lawmakers have gone as far as calling for establishing a dedicated agency to oversee AI development and deployment (Tarinelli, 2023).
Others, however, caution against reinventing the regulatory wheel, instead favoring the adaptation of existing laws (Corrigan and Daniels, 2024). While recent attention in AI policy is largely driven by generative AI, the need to regulate AI is not new. For years before ChatGPT was released to the public, regulators have applied sector-specific frameworks to address the use of predictive systems in employment, credit, housing, and public benefits (Consumer Financial Protection Bureau, 2022; Department of Housing and Urban Development, 2024; Jillson, 2021). Building on these efforts, those who regard generative AI as a “normal technology” support strengthening existing frameworks and institutional capacities to regulate generative applications rather than pursuing drastic policy interventions (Narayanan and Kapoor, 2025).
Although the debate remains unsettled on whether generative AI requires new laws, enforcing existing and future laws will require new tools. Over the past decade, policymakers have designed their regulatory toolkit around particular features of predictive AI. Many of these features, however, lose their relevance in a generative setting. In this paper, we identify four important differences between generative and predictive AI that policymakers should consider in governing both types of systems. We find that the training process of generative AI, its outputs, and its supply chain structure fundamentally depart from the assumptions that existing regulations were built on. These distinctions pose new challenges for policymaking, oversight, and accountability.
@@
Whether emerging technology drives policy has been a longstanding question technology law, including AI governance (Crootof and Ard, 2021; Calo, 2015; Jones, 2018). Drawing on past lessons, we offer three recommendations. First, that regulation should target use cases and impact, rather than the technical specification of AI systems. Second, that technical proxies should be used sparingly in regulation, with a careful balancing of risk, benefits, and a consideration of alternatives. Third, that regulation should focus on the entire risk chain for a use case and reinforce structural constraints that can limit harms, rather than focusing on AI systems alone.
2. Technical Distinctions
AI definitions shape-shift depending on the context in which the terms are used. Terms such as generative AI, general purpose AI, automated decision-making systems, and others are commonly used to describe systems that incorporate computational elements powered by machine learning or AI. Sometimes, the language refers to the underlying technology in the system, the nature of the end-use, or even the lack of a specific end-use.
@@ 2. Technical Distinctions
Our goal is to draw distinctions between models (and systems) built to perform specific tasks, typically referred to as predictive or discriminative, and those constructed without a specific task in mind, often referred to as generative (including those that are later fine-tined for a specific task). [^1] The EU AI Act, among others, makes this same distinction, but calls out the former systems as ‘AI’ and the latter as ‘general-purpose AI’. (Bommasani et al., 2021) refer to the former as ‘task-specific’ (and the latter as ‘task-agnostic’, but settle on ‘foundation’). We feel that the terms ‘predictive’ and ‘generative’ are useful for the distinctions that we wish to explore, but note that there are several different ways to refer to these systems without a clear consensus. In the following section, we highlight the key differences in the construction of these systems that shape their impact and should impact how policymakers govern their use.
2.1. Predictive Systems
@@
Predictive AI models are task-specific systems that are designed to predict a particular outcome and are almost exclusively built using supervised learning. Building a predictive model starts with a task described by training data consisting of input vectors and the “ground truth” predictions. The goal is to build a model that can correctly reproduce the ground truth predictions given the inputs, in the hope that this model will accurately predict outcomes for a new input.
@@
While this description is an oversimplification of model training, it captures the most relevant aspects of predictive AI for policymakers—there is an intended task, the task can be expressed in terms of specific training data, and the goal is to build a model that performs this task. (Selbst et al., 2019).
Predictive models are very sensitive to the exact nature of the task they are trained for, and how that task is manifested via the training data. There are notable examples where predictive systems fail to perform as expected because they were trained for tasks slightly different than their intended use (Obermeyer et al., 2019; Passi and Barocas, 2019).
2.2. Generative Systems
@@
When we turn to generative AI, the picture is much more complicated. Some descriptions draw distinctions from predictive models based on the nature of the output—content versus predictions (Organisation for Economic Co-operation and Development (2024), OECD). This is reflected in AI definitions in legislative proposals that insert the word ‘content’ into the list of outputs that might be produced by an AI system (Rhode Island State Legislature, 2025; Texas State Legislature, 2025; Iowa State Legislature, 2025).
@@
In this paper, we view a generative AI system as one that is built using self-supervision, that usually (though not necessarily) is described using billions of parameters, and that most importantly is not trained to perform any specific task beyond that which the self-supervision requires. [^2] To illustrate the difficulty even with this rendition, systems like BERT (Devlin et al., 2019) are built using self-supervision, but are not generative in the sense of producing new, fluent sequences of text. Self-supervision is perhaps the most distinctive feature of what characterizes generative AI models. In language models, for example, self-supervision proceeds by training the language model to fill in the blanks in lines of text in which words have been masked randomly. [^3] The model is trained using supervised learning with a specific task—predict the missing words—but using the training data itself as inputs and targets for training. This gives rise to the term “self-supervision.” For image models, the process is different, but is still self-supervised. The rationale for this form of training is that for a model to be able to correctly predict missing words in arbitrary sentences in text, it must be able to build an internal representation of meaning in text that will allow it to complete the sentences. Regardless of whether models are able to represent ‘meaning’, it is the case that models do end up constructing an internal representation space for words (or tokens) and use this internal representation to respond to prompts or other natural language queries. This internal representation is what allows generative AI systems to perform a wide range of downstream tasks for which it might not have ever been trained specifically.
@@
Most of the other elements used to define generative AI systems—the architecture of transformers, the apparent need for billions of parameters, and so on—are either consequences of the training process or reflections of prevailing design practices. These features represent our current understanding of how to build models effectively, but they are not necessary to capture the general-purpose task-agnostic nature of generative AI systems.
But what of the myriad further steps in training a generative AI model, like fine-tuning, reinforcement learning from human feedback, inference-tuning, and so on? A first point that we note is that most of these post-pretraining steps do not fundamentally change the fact that a generative model relies on an internal representation to perform a wide range of downstream tasks, even if they might customize the model to perform more effectively on certain tasks.
@@
Secondly, these downstream training steps apply to an already built model, rather than creating a new one from scratch. Unlike predictive models, where using the model for a new task usually involves rebuilding it from scratch, fine-tuning merely modifies an existing (and functioning) model, and can be performed without knowledge of how the original model was built. [^4] The knowledgeable reader will note that transfer learning is exactly the process of taking a given model and adapting it using small amounts of new data for a related task. While this is true, it is typically less common in predictive settings to employ transfer learning since it is more straightforward to rebuild the model. In turn, the builders of the base model need not have any knowledge of how fine-tuning might occur and for what purpose. A generative model that is fine-tuned to perform a prediction task is still a generative model for the purpose of the points we discuss in this paper.
3. What makes generative AI distinctive?
@@
In this section, we outline four key distinctions that set generative AI apart—its general-purpose design, challenges in evaluation, emergence of new legal concerns, and distributed value chain structure. These differences challenge the assumptions underlying existing regulatory instruments, revealing how oversight and accountability mechanisms designed for predictive systems lose their utility when applied to generative systems.
@@
3.1. Generative models make poor regulatory targets because of their generality and adaptability.
@@
Generative models are general-purpose and adaptable, able to handle unstructured inputs across many modalities (e.g., text, images, audio, etc.) As a result, the space of potential applications for generative models is vast. Through fine-tuning and prompting, the same foundation model can be adapted to perform remarkably different tasks without specific training. For example, GPT-4 can understand and generate human-like text across various topics and styles, process and generate images, engage in real-time verbal conversations, perform complex reasoning tasks, and write and debug code, all without fundamentally changing the underlying model (OpenAI, 2023).
@@
In contrast, predictive systems are designed for specific tasks. They streamline, support, and sometimes fully automate human decision-making processes. Although these systems vary in their technical implementation, they share a common function as tools for decision-making. This shared function gives rise to a shared set of risks across different applications. Whether applied to hiring, credit scoring, or tenant screening, predictive AI raises a similar set of concerns related to fairness, transparency, and accountability. This creates a coherent basis for regulators to establish overarching guardrails that target universal harms posed by predictive AI—one that draws on principles of due process, non-discrimination, and privacy. For example, the Blueprint for an AI Bill of Rights outlines common expectations for automated decision systems, such as the right to notice and explanation and protection from algorithmic bias, regardless of the domain (White House Office of Science and Technology Policy, 2022a). Similarly, the proposed AI Civil Rights Act seeks to prohibit the use, sale, or promotion of algorithmic decision-making systems that discriminate on the basis of protected characteristics (Markey, 2024).
@@
The generality and adaptability of generative AI breaks down this regulatory coherence. Generative systems can be put to a far more diverse set of uses and although these applications may stem from the same technical foundation, many will raise distinct risk profiles. Like predictive AI, generative AI can be used for decision-making, including in the high-stakes context of hiring (Gaebler et al., 2025), lending (Bowen III et al., 2024), and housing (Liu et al., 2024), thus invoking the same set of regulatory concerns. But it can also be used to generate children’s stories, which raises concerns over age-appropriate material and creative rights (UNICEF, 2021; U.S. Copyright Office, 2024), for drafting legal documents, where there may be concerns over hallucinations and the confidentiality risks of client information (The State Bar of California, 2023), and for coding assistance, which could introduce security vulnerabilities and lower code quality (Perry et al., 2023; Harding, 2025). While there are sector-specific frameworks that govern each of these use cases, there is no shared regulatory objective that connects them.
@@
Developing regulations that apply across such diverse applications would require abstracting away all of the details that actually account for the concerns raised by these specific applications. Because generative systems lack a shared purpose, the only features shared across applications tend to be technical properties of the models, such as compute usage and model size, which do not correspond to any particular applications of concern. As a result, when regulators attempt to regulate generative systems broadly, they are left with few footholds other than these technical properties. These features are attractive from a regulatory perspective because they are easy to define and measure. However, they are often poor proxies for the many harms that could arise from different applications.
@@
The result is a set of rules that are broad in scope, but shallow in substance. These rules are shallow because they only address the highest-level factors that could affect the risks posed by generative AI—typically only those observable at the model level without consideration of the deployment context. This creates a regulatory framework that appears expansive but lacks the specificity to ensure meaningful oversight. We elaborate on the limitations of treating generative models as the regulatory target in 4.1, where we instead recommend regulating the applications and impacts of these systems.
@@
3.2. Generative systems are difficult to evaluate due to their lack of task specificity.
@@
Evaluation of AI systems has become a touchstone for oversight. In many cases, evaluations are designed to assess system performance using quantitative metrics, such as accuracy and recall, that typically require a known ground truth for comparison. By systematically assessing AI system performance, evaluations provide enforcement mechanisms and remedies to contest false claims and ineffective systems (Raji et al., 2022; National Telecommunications and Information Administration, 2024). In other cases, evaluations are designed to assess models’ impacts. One common example is the use of evaluation to identify when predictive AI systems produce disparate impacts—that is, when they differentially benefit or burden different groups. From the E.U. AI Act to New York City local laws, policymakers have called for independent audits, impact assessments, and risk assessments in order to understand both basic functionality (i.e., performance) and harms (i.e., impacts), including discrimination, privacy, and physical safety (European Parliament, 2024; New York City Council, 2020; White House Office of Science and Technology Policy, 2022b). The widespread adoption of audit requirements and impact assessments in U.S. congressional bills (U.S. House of Representatives, 2023), state-level initiatives (California State Legislature, 2019; Colorado General Assembly, 2021), and international proposals (European Parliament, 2024; Innovation, Science and Economic Development Canada, 2023; Information Commissioner’s Office, 2023) reflects their importance in providing assurance and accountability in AI systems.
@@
Predictive systems are designed around tasks that map from specific input spaces to output spaces (see Section 2). This means developers are able to conduct evaluations within bounded problem domains where decision-making objectives can be clearly defined and measured. Typically, evaluators use a holdout set—data withheld from training that contains known outcomes that serve as the ground truth for assessing the quality of the models’ predictions (Donoho, 2017). By comparing the model’s outputs to these outcomes, evaluators can compute a set of metrics (e.g., accuracy) to measure how well the system performs on a predetermined task. The evaluation process naturally mirrors training by repurposing the same data pipelines and directly measuring the model’s performance on the task it was trained on (Kohavi, 1995).
@@
The task-specificity of predictive systems has three important implications for evaluations. First, evaluations of predictive systems have self-evident targets. Because the model is trained with a clear objective, evaluators can measure how well the system performs its intended function. Second, evaluations can be properly scoped. That is, evaluators can define the input space for each predictive system and bound its functionality to specific use cases. This allows evaluators to assess model performance, knowing that the model will only operate within this defined input space and purpose. Third, evaluations can be reproducible. When provided with the same inputs, outputs, and metrics, different evaluators can independently reproduce and verify the system’s performance. These three properties allow policymakers to establish certain expectations around the validity and reliability of evaluations. Although there remain technical and institutional challenges in evaluations for predictive systems, the expectations are that (i) evaluations should appropriately characterize the expected behavior of the systems relative to its defined objective (targeted), (ii) they should achieve sufficient coverage of the system’s intended uses (well-scoped), and (iii) the evaluation results should be consistent and repeatable across different independent evaluators (reproducible).
@@
The absence of task specificity in generative systems means that these properties fall away. Generative systems are not trained for specific tasks but rather with broader objectives to learn structural patterns in the data (e.g. predicting the next word given the previous words). As such, there is no clear target against which to evaluate their performance (Wallach et al., 2024). Evaluators instead target general capabilities or properties, such as linguistic understanding (Wang et al., 2020), reasoning (Chollet, 2019; Cho et al., 2023), and biases (Parrish et al., 2022; Birhane et al., 2021; Talat et al., 2022). These capabilities are characterized by the model’s responses to a set of custom-built and annotated scenarios, typically in the form of multiple-choice or binary questions (et al., 2023b, a; Hendrycks et al., 2021; Wang et al., 2020). A scenario is meant to instantiate a desired use case for generative models, and a collection of scenarios form benchmarks that guide and measure progress in AI development (et al., 2023b).
@@
While capabilities-based evaluations are meant to produce evidence of likely performance on a broad range of possible tasks that would presumably draw on relevant capabilities, the relationship between capabilities and performance on specific, real-world tasks is often quite tenuous. As many have observed, the particular tasks used in capabilities-focused benchmarks don’t often resemble the tasks that generative models are used to perform in practice (Subramonian et al., 2023; Raji et al., 2021; McIntosh et al., 2024; Davis, 2023; Bowman and Dahl, 2021; Ethayarajh and Jurafsky, 2021; McCoy et al., 2019; Ribeiro et al., 2020).

### v3

# Distinguishing Task-Specific and General-Purpose AI in Regulation

@@ Abstract.
Over the past decade, policymakers have developed a set of regulatory tools to ensure AI development aligns with key societal goals. Many of these tools were initially developed in response to concerns with task-specific AI and therefore encode certain assumptions about the nature of AI systems and the utility of certain regulatory approaches. With the advent of general-purpose AI (GPAI), however, some of these assumptions no longer hold, even as policymakers attempt to maintain a single regulatory target that covers both types of AI.
In this paper, we identify four distinct aspects of GPAI that call for meaningfully different policy responses. These are the generality and adaptability of GPAI that make it a poor regulatory target, the difficulty of designing effective evaluations, new legal concerns that change the ecosystem of stakeholders and sources of expertise, and the distributed structure of the GPAI value chain.
In light of these distinctions, policymakers will need to evaluate where the past decade of policy work remains relevant and where new policies, designed to address the unique risks posed by GPAI, are necessary. We outline three recommendations for policymakers to more effectively identify regulatory targets and leverage constraints across the broader ecosystem to govern GPAI.
AI Governance, Regulatory Target, AI Value Chain
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Symposium on Computer Science and Law; March 03–05, 2026; Berkeley, CA, USAbooktitle: Symposium on Computer Science and Law (CSLAW ’26), March 03–05, 2026, Berkeley, CA, USAdoi: 10.1145/3788646.3789523isbn: 979-8-4007-2447-3/2026/03ccs: Social and professional topics Governmental regulations
@@ 1. Introduction
Recent advances in AI have drawn intense interest from policymakers. In the past year alone, policymakers have proposed and enacted a litany of governance measures, from the European Union AI Act to a growing number of state and local initiatives in the United States (European Parliament, 2024; Curry, 2024). This wave of legislative activity has generated new confusion and contestation over how AI should be defined and governed. Some treat generative and general-purpose AI (GPAI) as a fundamentally new class of technology whose scale and autonomy distinguish it from others (Bostrom, 2014; Altman et al., 2023), advocating for entirely new laws and institutions (Nichols, 2023; Bureau of Industry and Security, U.S. Department of Commerce, 2025). Others caution against reinventing the regulatory wheel, instead favoring the adaptation of existing laws (Corrigan and Daniels, 2024).
A notable source of this confusion is the failure to distinguish between task-specific systems and GPAI. Some regulatory efforts draw this line explicitly – for instance, the E.U. AI Act imposes distinct obligations on GPAI providers (Boine and Rolnick, 2023). Many state and local proposals, however, do not. New York’s A6453-B, for example, defines AI as a system that “can … make predictions, recommendations, or decisions,” yet its provisions targeting model-weight theft are clearly aimed at generative systems (New York State Assembly, 2025). We also see the reverse: laws designed around task-specific systems being imposed on GPAI. New Mexico HB 60 describes AI as a technology that can be used to “generate outputs, including content, decisions, predictions, or recommendations,” yet it requires the disclosure of “sources or owners of the datasets” and the “factors in the data” used to produce outputs — assumptions that fit task-specific models with identifiable datasets and features, but not GPAI models (Chandler et al., 2025).
This confusion matters because, over the past decade, policymakers have designed their regulatory toolkit around particular features of task-specific AI, but many of these features lose their relevance in a generative setting. For years before ChatGPT was released to the public, regulators have applied sector-specific frameworks to address the use of task-specific systems in employment, credit, housing, and public benefits (Bureau, 2022; Department of Housing and Urban Development, 2024; Jillson, 2021). These regimes were built around assumptions of how these systems are trained and evaluated for well-defined tasks, and how they are deployed to specific domains with legal guardrails. GPAI, however, breaks many of these assumptions.
In this paper, we identify four important differences between general-purpose and task-specific AI that policymakers should consider in governing both types of systems. First, the generality and adaptability of GPAI makes it a poor regulatory target: a single model can be repurposed across many domains and use cases, so rules aimed at the model itself may fail to track where risk actually arises. Second, GPAI systems are difficult to evaluate in reliable ways because they lack task specificity. Third, GPAI raises new legal concerns that were absent or far less salient in task-specific settings, shifting the ecosystem of stakeholders and expertise required for oversight. Fourth, GPAI is developed and deployed through a highly distributed value chain which complicates the attribution of responsibility. These distinctions create governance challenges that necessitate new regulatory tools to enforce both existing and new laws.
@@
In light of these distinctions, policymakers will need to evaluate where the past decade of policy work remains fit for purpose and where new policies are needed to address the risks unique to generative AI. Consequently, we offer three recommendations. First, that regulation should target use cases and impact, rather than the technical specification of AI systems. Second, that technical proxies should be used sparingly in regulation, with a careful balancing of risk, benefits, and a consideration of alternatives. Third, that regulation should focus on the entire risk chain for a use case and reinforce structural constraints that can limit harms, rather than focusing on AI systems alone.
2. Conceptual Distinctions
@@
In this paper, we draw a distinction between models built to perform specific tasks, typically referred to as task-specific, predictive, or discriminative, and those constructed without a specific task in mind, often referred to as general-purpose (including those that are later fine-tuned for a specific task) or generative. While the distinction we raise may not capture every technical detail, it reflects the features most pertinent to policymakers. [^1] The EU AI Act, among others, makes this same distinction, but calls out the former systems as ‘AI’ and the latter as ‘general-purpose AI’. Bommasani et al. (2021) refer to the former as ‘task-specific’ (and the latter as ‘task-agnostic’, but settle on ‘foundation’). We feel that the terms ‘task-specific’ and ‘general-purpose’ are useful for the distinctions that we wish to explore, but note that there are several different ways to refer to these systems, and there is no clear consensus.
2.1. Task-specific AI systems
@@
Task-specific AI systems are designed to predict a particular outcome and are almost exclusively built using supervised learning. Building such a system starts with a task described by training data consisting of input vectors and the “ground truth” predictions. The goal is to build a system that can correctly reproduce the ground truth predictions given the inputs, in the hope that it will accurately predict outcomes for a new input.
@@
While this description is an oversimplification of model training, it captures the most salient aspects of task-specific AI for policymakers—there is an intended task, the task can be expressed in terms of specific training data, and the goal is to build a system that performs this task. (Selbst et al., 2019).
2.2. General-purpose AI systems
@@
When we turn to GPAI, the picture is much more complicated. Some policy definitions draw distinctions from task-specific systems based on the nature of the output—content versus predictions (Organisation for Economic Co-operation and Development (OECD), 2024). This is reflected in AI definitions in legislative proposals that insert the word ‘content’ into the list of outputs that might be produced by an AI system (Rhode Island State Legislature, 2025; Texas State Legislature, 2025; Iowa State Legislature, 2025).
@@
In line with these policy discussions, we view a GPAI system as one that is built using self-supervision, that usually (though not necessarily) is described using billions of parameters, and that most importantly is not trained to perform any specific task beyond that which the self-supervision requires. [^2] To illustrate the difficulty even with this rendition, systems like BERT (Devlin et al., 2019) are built using self-supervision, but are not generative in the sense of producing new, fluent sequences of text. Self-supervision is perhaps the most distinctive feature of what characterizes GPAI models. In language models, for example, self-supervision proceeds by training the language model to fill in the blanks in lines of text in which words have been masked randomly. [^3] The model is trained using supervised learning with a specific task—predict the missing words—but using the training data itself as inputs and targets for training. This gives rise to the term “self-supervision.” For image models, the process is different, but is still self-supervised. The rationale for this form of training is that for a model to be able to correctly predict missing words in arbitrary sentences in text, it must be able to build an internal representation of meaning in text that will allow it to complete the sentences. Regardless of whether models are able to represent ‘meaning’, it is the case that models do end up constructing an internal representation space for words (or tokens) and use this internal representation to respond to prompts or other natural language queries. This internal representation is what allows GPAI systems to perform a wide range of downstream tasks for which it might not have ever been trained specifically. [^4] While some self-supervised systems are built for specific tasks, policy discussions focus on GPAI systems where self-supervision enables broad, task-agnostic capabilities.
@@
Most of the other elements used to define GPAI systems—the architecture of transformers, the apparent need for billions of parameters, and so on—are either consequences of the training process or reflections of prevailing design practices. These features represent our current understanding of how to build models effectively, but they are not necessary to capture the task-agnostic nature of GPAI systems.
@@
But what of the myriad further steps in training a GPAI model, like fine-tuning, reinforcement learning from human feedback, inference-tuning, and so on? First, most of these post-pretraining steps do not fundamentally change the fact that a GPAI model relies on an internal representation to perform a wide range of downstream tasks, even if they might customize the model to perform more effectively on certain tasks. Second, these downstream training steps apply to an already built model, rather than creating a new one from scratch. Unlike task-specific models, where using the model for a new task usually involves rebuilding it from scratch, fine-tuning merely modifies an existing (and functioning) model, and can be performed without knowledge of how the original model was built. [^5] The knowledgeable reader will note that transfer learning is exactly the process of taking a given model and adapting it using small amounts of new data for a related task. While this is true, it is typically less common in predictive settings to employ transfer learning since it is more straightforward to rebuild the model. In turn, the builders of the base model need not have any knowledge of how fine-tuning might occur and for what purpose. A GPAI model that is fine-tuned to perform a prediction task is still a GPAI model for the purpose of the points we discuss in this paper.
3. What makes general-purpose AI distinctive?
@@
In this section, we outline four key distinctions that set GPAI apart—its general-purpose design, challenges in evaluation, emergence of new legal concerns, and distributed value chain structure. These differences challenge the assumptions underlying existing regulatory instruments, revealing how oversight and accountability mechanisms designed for task-specific systems lose their utility when applied to GPAI systems.
@@
3.1. General-purpose AI systems make poor regulatory targets because of their generality and adaptability.
@@
GPAI systems are general-purpose and adaptable, able to handle unstructured inputs across many modalities (e.g., text, images, audio, etc.). As a result, the space of potential applications for GPAI is vast. Through fine-tuning and prompting, the same base model can be adapted to perform remarkably different tasks without specific training. For example, GPT-4 can understand and generate human-like text across various topics and styles, process and generate images, engage in real-time verbal conversations, perform complex reasoning tasks, and write and debug code, all without fundamentally changing the underlying model (OpenAI, 2024b).
@@
In contrast, task-specific systems are purpose-built to perform a defined task. They streamline, support, and sometimes fully automate human decision-making processes. Although these systems vary in their technical implementation, they share a common function as tools for decision-making. This shared function gives rise to a shared set of risks across different applications. Whether applied to hiring, credit scoring, or tenant screening, task-specific AI raises a similar set of concerns related to fairness, transparency, and accountability. This creates a coherent basis for regulators to establish overarching guardrails that target universal harms posed by task-specific AI—one that draws on principles of due process, non-discrimination, and privacy. For example, the Blueprint for an AI Bill of Rights outlines common expectations for automated decision systems, such as the right to notice and explanation and protection from algorithmic bias, regardless of the domain (White House Office of Science and Technology Policy, 2022a). Similarly, the proposed AI Civil Rights Act seeks to prohibit the use, sale, or promotion of algorithmic decision-making systems that discriminate on the basis of protected characteristics (Markey, 2024). The comprehensive AI regulation passed in Colorado in 2024 similarly sets requirements for developers of ”high-risk” AI systems across a wide range of domains (Colorado General Assembly, 2024).
@@
The generality and adaptability of GPAI breaks down this regulatory coherence. GPAI systems can be put to a far more diverse set of uses and although these applications may stem from the same technical foundation, many will raise distinct risk profiles. Like task-specific AI, GPAI can be used for decision-making, including in the high-stakes context of hiring (Gaebler et al., 2025), lending (Bowen III et al., 2024), and housing (Liu et al., 2024), thus invoking the same set of regulatory concerns. But it can also be used to generate children’s stories, which raises concerns over age-appropriate material and creative rights (UNICEF, 2021; U.S. Copyright Office, 2024), for drafting legal documents, where there may be concerns over hallucinations and the confidentiality risks of client information (The State Bar of California, 2023), and for coding assistance, which could introduce security vulnerabilities and lower code quality (Perry et al., 2023; Harding, 2025). While this generality and adaptabiltiy is not inherently problematic, as there are sector-specific frameworks that govern each of these use cases, the challenge arises when treating GPAI as the regulatory target (i.e., the specific entity that regulation is designed to govern).
@@
Because there is no shared regulatory objective (i.e., the specific goal or intended outcome that regulation aims to achieve) across generative use cases, efforts to develop regulations at the model level may be misguided. Model-level regulations that apply across diverse applications would require abstracting away all of the details that actually account for the concerns raised by these specific applications. Because GPAI systems lack a shared purpose, the only features shared across applications tend to be technical properties of the models, such as compute usage and model size, which do not correspond to any particular applications of concern. As a result, when regulators attempt to pursue broad upstream controls specifically targeting GPAI, they are left with few footholds other than these technical properties. These features are attractive from a regulatory perspective because they are easy to define and measure. However, they are often poor proxies for the many harms that could arise from different applications.
@@
Thus, regulation targeting GPAI systems often yields rules that are broad in scope, but shallow in substance. These rules are shallow because they only address the highest-level factors that could affect the risks posed by GPAI—typically only those observable at the model level without consideration of the deployment context. This creates a regulatory framework that appears expansive but lacks the specificity to ensure meaningful oversight. We elaborate on the limitations of regulating at the model level in 4.1, where we instead recommend regulating the applications and impacts of these systems.
@@
3.2. General-purpose AI systems are difficult to evaluate due to their lack of task specificity.
@@
Evaluation of AI systems has become a touchstone for oversight. In many cases, evaluations are designed to assess system performance using quantitative metrics, such as accuracy and recall, that typically require a known ground truth for comparison. By systematically assessing AI system performance, evaluations provide enforcement mechanisms and remedies to contest false claims and ineffective systems (Raji et al., 2022; National Telecommunications and Information Administration, 2024). They are crucial to determining whether a system is fit for purpose and sufficiently reliable for deployment. In other cases, evaluations are designed to assess models’ impacts. One common example is the use of evaluation to identify when task-specific AI systems produce disparate impacts—that is, when they differentially benefit or burden different groups. From the E.U. AI Act to New York City local laws, policymakers have called for independent audits, impact assessments, and risk assessments in order to understand both basic functionality (i.e., performance) and harms (i.e., impacts), including discrimination, privacy, and physical safety (European Parliament, 2024; New York City Council, 2020; White House Office of Science and Technology Policy, 2022b). The widespread adoption of audit requirements and impact assessments in U.S. congressional bills (U.S. House of Representatives, 2023), state-level initiatives (California State Legislature, 2019; Colorado General Assembly, 2021), and international proposals (European Parliament, 2024; Innovation, Science and Economic Development Canada, 2023; Information Commissioner’s Office, 2023) reflects their importance in providing assurance and accountability in AI systems.
@@
Task-specific systems are designed around tasks that map from specific input spaces to output spaces (see Section 2). This means developers are able to conduct evaluations within bounded problem domains where decision-making objectives can be clearly defined and measured. Typically, evaluators use a holdout set—data withheld from training that contains known outcomes that serve as the ground truth for assessing the quality of the models’ predictions (Donoho, 2017). By comparing the model’s outputs to these outcomes, evaluators can compute a set of metrics (e.g., accuracy) to measure how well the system performs on a predetermined task. The evaluation process naturally mirrors training by repurposing the same data pipelines and directly measuring the model’s performance on the task it was trained on (Kohavi, 1995).
@@
The task-specificity of these systems has three important implications for evaluations. First, evaluations of task-specific systems have self-evident targets. Because the model is trained with a clear objective, evaluators can measure how well the system performs its intended function. Second, evaluations can be properly scoped. That is, evaluators can define the input space for each task-specific system and bound its functionality to specific use cases. This allows evaluators to assess model performance, knowing that the model will only operate within this defined input space and purpose. Third, evaluations can be reproducible. When provided with the same inputs, outputs, and metrics, different evaluators can independently reproduce and verify the system’s performance. These three properties allow policymakers to establish certain expectations around the validity and reliability of evaluations. Although there remain technical and institutional challenges in evaluations for task-specific systems, the expectations are that (i) evaluations should appropriately characterize the expected behavior of the systems relative to its defined objective (targeted), (ii) they should achieve sufficient coverage of the system’s intended uses (well-scoped), and (iii) the evaluation results should be consistent and repeatable across different independent evaluators (reproducible).
@@
The absence of task specificity in GPAI systems means that these properties fall away. GPAI systems are not trained for specific tasks but rather with broader objectives to learn structural patterns in the data (e.g. predicting the next word given the previous words). As such, there is no clear target against which to evaluate their performance (Wallach et al., 2024). Evaluators instead target general capabilities or properties, such as linguistic understanding (Wang et al., 2020), reasoning (Chollet, 2019; Cho et al., 2023), and biases (Parrish et al., 2022; Birhane et al., 2021; Talat et al., 2022). These capabilities or properties are characterized by the model’s responses to a set of custom-built and annotated scenarios, typically in the form of multiple-choice or binary questions (et al., 2023b, a; Hendrycks et al., 2021; Wang et al., 2020). A scenario is meant to instantiate a desired use case for GPAI models, and a collection of scenarios form benchmarks that guide and measure progress in AI development (et al., 2023b).
@@
While capabilities-based evaluations are meant to produce evidence of likely performance on a broad range of possible tasks that would presumably draw on relevant capabilities, the relationship between capabilities and performance on specific, real-world tasks is often quite tenuous. As many have observed, the particular tasks used in capabilities-focused benchmarks don’t often resemble the tasks that GPAI models are used to perform in practice (Subramonian et al., 2023; Raji et al., 2021; McIntosh et al., 2024; Davis, 2023; Bowman and Dahl, 2021; Ethayarajh and Jurafsky, 2021; McCoy et al., 2019; Ribeiro et al., 2020).